Hello,
Please help me to solve a vector problem.
Given the vectors:
a=i+2y-k
b=2i-y+k
c=i+2k
Find vectors:
a+2b-c
a*c
c·a
(a*b)·c
Thank you very much!

Answer by Alan3354(69443)

Given the vectors:
a=i+2y-k
b=2i-y+k
c=i+2k
Find vectors:
a+2b-c
---------------
Is it i, y & k?
---

2b = 4i - 2y + 2k
 a =  i + 2y -  k
-c = -i      - 2k
----------------------------- Add
     4i - k
==================

a*c
If you mean the cross-product:
 i    y    k
 1    2   -1
-1    0   -2
----
= i*(-4) - y*(-2-1) + k*(2)
= -4i + 3y + 2k

======================
c·a
a=i+2y-k
c=i+0y+2k
--> 1*1 -1*2 = -1
=====================
(a*b)·c
----
Do like the ones above.